OpenStudy (anonymous):

Solve: 2a - b + c = 0 a - b = -3 2a + b = -3

OpenStudy (wolf1728):

Add these 2 equations: a - b = -3 2a + b = -3 equals 3a = -6 a = -2 b= 1 Substitute these values into this equation to get c: 2a - b + c = 0